A lady whose mass is 50 kg stands on a scale in an elevator. As the elevator approaches the ground ?oor from above, it is slowing at a rate of 3 m/s2. What is the reading on the scale for the lady
Answer #1

let the normal reaction on lady be N (this is reading on scale as scale does not measure weight but gives the value of normal reaction)

so applying newton's second law

N - mg = ma

N = mg+ ma

N = 50*9.8 + 50*3 = 640 N

reason a is positive is i have taken y as positive direction

since elevator is slowing it value is negative in downward direction